OpenStudy (anonymous):

Easy medal. In which expression should the exponents be multiplied? (5^3)^-3 (1/3)^4 • (1/3)^7 3^7/3^13 4^5 • 4^2

Nnesha (nnesha):

exponent rules \[\huge\rm (x^m)^n=x^{m \times n}\]\[\huge\rm x^m \times x^n =x^{m +n}\] \[\huge\rm \frac{ x^m }{ x^n }= x ^{m-n}\]

Nnesha (nnesha):

thats all you need to find an answer.

OpenStudy (anonymous):

Which one of those should be multiplied though

OpenStudy (anonymous):

Would it be b or d?

Nnesha (nnesha):

when you multiply same bases you should add their exponents exponent rule number 2!